Semiautomatic process has a fixed cost of $40,000 per year and variable cost of $30 per unit. An automatic process has fixed cost of $88,000 per year and a variable cost of $22 per unit. At what production level per year will the two alternatives break even?
Answer #1

Answer

The breakeven quantity is at TC of one is equal to other

TC=FC+VC

For Semiautomatic Process

FC=40000

VC=30 per unit

TC=40000+30Q

For Automatic process

FC=88000

VC=22 per unit

TC=88000+22Q

TC(semiautomatic)=TC(automatic)

40000+30Q=88000+22Q

8Q=48000

Q=6000

The two alternatives will break even at a production level of 6000 per year.